# Assessing the dynamical assumptions in Tsirelson inequality tests of non-classicality in harmonic oscillators

### v1

@@ III.2 Method 1: Macrorealistic Description of Sign Changes
Step 1: The full data set comprises all three correlators $C_{12},C_{23},C_{13})$, measured non-non-invasively, and $\langle A\rangle$ via a Tsirelson test.
@@
Step 3: Finally, we ask the question: is it possible that a non-zero no-crossing probability (that breaks uniform precession) explains the full Tsirelson violation? The first term on the right-hand side of Eq. 3.2 is positive, and we have calculated $\frac{1}{2}(1+\langle A\rangle)=-0.0598$ and $L_{1}=0.0421$ (from Appendix F). It is immediately clear that the violation of uniform precession is not sufficient to explain the Tsirelson inequality violation; hence non-zero quantum interferences are present. This is a crucial result, showing that this method has proven successful in identifying that uniform precession is satisfied well enough that quantum behaviour must be implied.
We also note that Eq. 3.2 can instead be written in terms of the two-time quasi-probabilities (Eq. 1.6), with no interferences:
@@
From Eq. 3.6 and Eq. 3.7, we see that the Tsirelson violation, $\frac{1}{2}(1+\langle A\rangle)$, will itself respect the Lüders bound if uniform precession is satisfied exactly ($L_{1}=0$), but in general can exceed this bound, as is the case with the maximally violating state in the full infinite-dimensional QHO Hilbert space, which Tsirelson found to be $\frac{1}{2}(1+\langle A\rangle)=-0.13<-0.125$ [1].
@@
with the appropriate normalisation factor $N$, yields a Tsirelson violation of $\langle A\rangle=-1.0795<-1$. Thus, the method of assessing UP with an LG quantity can also be used upon such states, which may be of interest experimentally, for Gaussian state superpositions can be easier to create in a laboratory setting [71].
@@ III.3 Method 2: Three-time Sequential Measurements
Here $I$ is the interference term, and its values for each of the four cases are given below in Table 2. Both classically and quantum-mechanically, the triple correlator satisfies $|D|\leq 1$, which makes the first term $\frac{1}{2}(1-D)$ manifestly non-negative, so that Eq. 3.9 is the explicit form of Eq. 3.1 in this method.
Step 3: We compare the size of the UP violation, $2|\Delta p|$, with the size of the Tsirelson violation $\frac{1}{2}|(1+\langle A\rangle)|$, noting once again that the interference terms would vanish in a purely classical model of the violation.
Table 2 catalogues our numerical results with this approach, considering each of the four aforementioned distributions. We compare the extent of a lower-bound violation that would be possible from only the breaking of the dynamical assumption in a classical model with the true value of the Tsirelson violation, using the same state as before (Eq. 2.10). With the three-time sequential measurement probability $\Delta p_{123}$ as the chosen measure of UP, we calculate that $2|\Delta p_{123}|=0.0191<0.0598$, showing once again that the breaking of the dynamical assumption falls significantly short of accounting for the full extent of the Tsirelson violation, thereby implying the presence of quantum-mechanical interference term.
When weak measurements are employed, UP is satisfied only marginally in the case of $\Delta p_{1(2)3}$; and when the quasi-probability distribution $\Delta q$ is instead used, which produces no interference terms, the UP violation is exactly commensurate with the Tsirelson violation. This is to be expected, since $\Delta q=0$ is a sufficient condition to derive the Tsirelson inequality, which follows directly from the moment expansion form of Eq. 1.7. The quasi-probabilities, though they can be negative, respect the sum rules for probabilities; when a measurement is rendered weak, the probability sum rule holds valid over that particular measurement time, as in $p_{1(2)3}(+,+,+)+p_{1(2)3}(+,-,+)=p_{13}(+,+)$.
@@
An alluring question is whether it is possible, by some variation of the parameters for the quantum state, to render $\Delta p_{123}=0$ exactly, whilst remaining within a Tsirelson violating region (such as that depicted in Fig. 4), which may serve the poetic justice of “total satisfaction” of the dynamical assumption, whereby the Tsirelson violation is solely due to effects quantum-mechanical in origin. In fact—unlike in Method 1, wherein setting $L_{1}=0$ leads trivially to the satisfaction of the Tsirelson inequality—it is possible in this case to set $\Delta p_{123}=0$. We find by using a constrained optimisation procedure via Lagrange multipliers that a simple state, still within the $N=6$ subspace, albeit now with complex coefficients, can be found to produce $\langle A\rangle=1.0756$ with $\Delta p_{123}=0$. This procedure is documented in Appendix J.
@@ IV.1 Dwell time
The dwell time is a measure of the fraction of time the particle spends in $x>0$. Classically, this value should be $1/2$ over a full period ($\tau=2\pi$).
This can readily be turned into a self-adjoint operator, $\hat{T}_{D}(\tau)$ [72], and we can take its expectation value, $\langle\hat{T}_{D}(\tau)\rangle$. This quantum-mechanical dwell time can be ascertained experimentally by measuring $\langle\hat{\theta}(x(t))\rangle$ at multiple times.
@@
We see that the dwell time, integrated over a full period, is exactly the classical value. [^2] In fact, the dwell time operator in the energy eigenbasis is proportional to the identity. By this measure, a classical and quantum oscillator are indistinguishable.
@@
where $\tau$ is some time on the interval $0,2\pi]$, as the result was obtained using the mean value theorem. Thus, we see that the dependence of $\langle A\rangle$ upon the average dwell time drops out, since $\langle\hat{T}(2\pi)\rangle$ assumes the classical value, and the Tsirelson violation is fundamentally rooted in the probability current at the origin.
@@ IV.2 Crossing number
Another simple property of classical uniform precession, which is closely linked to Tsirelson’s dynamical assumption, is that trajectories cross the origin exactly twice over a period (barring the singular case of a particle sitting at rest at the phase-space origin). Formally, we can compute crossings over a time interval $\tau$ using
@@
We observe an asymptotic tendency towards the classical value of two crossings in the limit of high energy eigenstates, which is depicted in Fig. 6. Moreover, we calculate for the Zaw–Scarani state an average crossing number of $\langle\hat{N}_{c}(2\pi)\rangle_{Z}\approx 1.49$, with a standard deviation of $\Delta\hat{N}_{c}(2\pi)_{Z}\approx 0.2$. Over a third of a period, between measurement times, $\hat{N}_{c}$ remains diagonal for states in $\mathcal{H}_{0}$, and so $\langle\hat{N}_{c}(2\pi/3)\rangle_{Z}\approx 0.50$ with a standard deviation of $\Delta\hat{N}_{c}(2\pi)_{Z}\approx 0.07$. This means that the probabilities $p(+,+,+)$ and p(,,)p(-,-,-), which underpin the assumption of uniform precession in the Tsirelson scheme, must be very small indeed, since this would require either no crossings at all or at least two crossings between any pair of measurement times. The spread in crossings over a third of a period suggests that this situation is statistically very unlikely; nonetheless, we have not found an exact connection to the relevant probabilities.
@@ IV.3 Implications for trajectory models
We have seen that the quantum-mechanical properties of the dwell time are identical to its classical properties and that as for the crossing number, its quantum-mechanical properties are qualitatively the same as classical and quantitatively similar. These features suggest that measurements of the dwell time and crossing number may be used to give indirect but useful bounds on the size of uniform precession violations. These quantities are, in principle, related to the three-time probabilities used in Section III, but any such relation will be non-unique, and we have not been able to find a particularly useful one. Therefore, we proceed instead using a simpler, more heuristic analysis.
@@
The properties of the dwell time, Eq. 4.3, over the time interval $0,2\pi]$ indicate that the underlying trajectories must each spend equal amounts of time in $x>0$ and $x<0$. However, as noted, the dwell time must have the same properties over any shifted time interval of the same size. The implication in phase space is the trajectories must spend equal times above or below any infinite radial line. Thus, the set of trajectories must have rotational symmetry, which suggests that the only possibility is a set of circles around the origin with constant angular velocity along each circle. This still allows a set of circular trajectories with a spread of possible constant angular velocities, which will cause a violation of uniform precession. However, trajectories with angular velocities faster or slower than the classical one for the given Hamiltonian will create non-classical values for the crossing number. But as the crossing number results above show, and as noted, the probabilities for four or more crossings or no crossings, which would be responsible for violations of uniform precession, are in fact very small.
@@ IV.4 Classical versus quantum dynamics
Finally, we make some general remarks about dynamics in classical versus quantum theories and why we might expect, on general grounds, that the violation of uniform precession be small. The predictions of quantum theory are very different to the predictions of classical physics, and most of these differences trace back to the simple fact that quantum theory allows a much richer variety of states than does classical physics—for example, through Wigner functions that can be negative versus classical phase-space probabilities. This is responsible for some of the most striking quantum phenomena, such as Bell inequality violations.
However, classical and quantum physics look at lot more similar at the level of dynamics, and there are many results in the history of quantum theory that prove this. For example, wave packets—or more generally, phase-space localised quasi-projectors—evolve to a good approximation along classical paths (see for example Ref. [73]). This in turn is related to the fact that the Wigner function, for a class of potentials, has approximately classical evolution. This behaviour can be affected by monitoring (e.g. with projection operators onto ranges of position at intermittent times), but this effect is not large unless the projections are onto very small regions and the projectors are applied very frequently, which is determined by the energy time $\hbar/\langle E\rangle$ [74]. (For the Zaw–Scarani state considered in this paper, the energy time is calculated to be $1/12$ of a period.) Furthermore, specific models, such as LG2 violations with a Gaussian initial state, produce some non-classicality, but only about $30\%$ of the level of non-classicality that can be produced by a negative Wigner function [64, 65]. Of course, this discussion involves quantum-mechanical properties. However, the sceptical macrorealist could confirm these ideas using control experiments in which the dynamics of a system in a set of initially classical-like states is measured.
@@ V Summary and Discussion
But, as we noted in Section I, there is a crucial loophole in the Tsirelson inequality: particularly, that violations could ensue from the failure of the dynamical assumption of uniform precession in a classical system, meaning that a Tsirelson violation is not a sufficient condition for genuine quantum behaviour. Therefore, in this paper we have developed a subsidiary protocol to disentangle violations of macrorealism from the breaking of the dynamical assumption—to which end we have embedded the Tsirelson inequality within the wider Leggett–Garg formalism.
@@
In Section IV, we presented further qualitative evidence that uniform precession is satisfied well in states that violate the Tsirelson inequality, by considering the heuristic notions of dwell times and the crossing number operator. We found that the quantum-mechanical dwell time exactly obeys the classical bound, and that the total number of origin crossings is approximately the classical value. On their own, these stand as interesting results, further elucidating the various ways in which uniform precession is and is not equivalent in the classical and quantum cases. But we also further interpreted these quantities in the context of underlying trajectory models, noting that they greatly restrict the extent to which underlying trajectories can be erratic. We also came upon the global property that Tsirelson violations are linked to the probability currents through the origin, whereby weakly-flowing states, or the classicalised currents of quantum states, never produce a violation.
@@ Acknowledgements.
We are grateful to Clement Mawby for useful discussions and Lin Htoo Zaw for bringing to our attention a result regarding superpositions of coherent states.
@@ Appendix A Tsirelson Quantity from the Wigner Function
The Wigner quasiprobability distribution $W_{\rho}(x,p)$ is given by
$W_{\rho}(x,p)=\frac{1}{2\pi i}\text{Tr}\left(\rho\ e^{\frac{i\pi}{2}\left[(\hat{X}-x)^{2}+(\hat{P}-p)^{2}\right]}\right)\ .$ (A1)
Thus, the Tsirelson quantity can also be written as
@@
where $W_{A}(x,p)$ is the phase-space representation of the operator $A$, obtained from its Weyl transform:
@@ E.1 Proof Strategy
    The largest matrix element, in magnitude, on a diagonal i.e for fixed $m$, $Q_{n,n+2m+1}$ is $Q_{0,2m+1}$.
@@
    The non-zero matrix elements decrease in magnitude across the 0th row i.e $Q_{0,2m+1}$ is a decreasing function in $m$.
@@ Appendix F Calculation of correlators in the quantum harmonic oscillator
$e^{i\hat{H}\tau}=\sum_{m=0}^{\infty}e^{iE_{m}\tau}|m\rangle\langle m|\ ,$ (F5)
@@
$C_{12}=\frac{1}{2}\sum_{n,k=0}^{N}c_{n}^{*}c_{k}\langle n|\hat{Q}_{1}\sum_{m=0}^{\infty}e^{iE_{m}\tau}|m\rangle\langle m|\hat{Q}_{1}e^{-iE_{k}\tau}|k\rangle+\frac{1}{2}\sum_{n,k=0}^{N}c_{n}^{*}c_{k}\langle n|e^{iE_{n}\tau}\hat{Q}_{1}\sum_{m=0}^{\infty}e^{-iE_{m}\tau}|m\rangle\langle m|\hat{Q}_{1}|k\rangle\ ,$ (F6)
@@
$\frac{1}{2}\sum_{n,k=0}^{N}\sum_{m=0}^{\infty}e^{i(E_{m}-E_{k})\tau}c_{n}^{*}c_{k}Q_{nm}Q_{mk}+\frac{1}{2}\sum_{n,k=0}^{N}\sum_{m=0}^{\infty}e^{i(E_{n}-E_{m})\tau}c_{n}^{*}c_{k}Q_{nm}Q_{mk}\ .$ (F7)
@@
$\frac{1}{2}\sum_{n,k=0}^{N}\sum_{m=0}^{\infty}e^{i(E_{m}-E_{k})\tau}c_{n}^{*}c_{k}Q_{nm}Q_{mk}+\frac{1}{2}[\sum_{n,k=0}^{N}\sum_{m=0}^{\infty}e^{i(E_{m}-E_{k})\tau}c_{n}^{*}c_{k}Q_{nm}Q_{mk}]^{*}\ .$ (F8)
@@
$\mathrm{Re}\{\sum_{n,k=0}^{N}\sum_{m=0}^{\infty}e^{i(E_{m}-E_{k})\tau}c_{n}^{*}c_{k}Q_{nm}Q_{mk}\}\ .$ (F9)
Taking $E_{n}=n\omega$ (with $\hbar=1$), ignoring the vacuum contribution $\frac{1}{2}\hbar\omega$, we obtain

### v2

@@ III.2 Method 1: Macrorealistic Description of Sign Changes
Step 1: The full data set comprises all three correlators, $C_{12},C_{23},C_{13}$, measured non-invasively, and $\langle A\rangle$ via a Tsirelson test.
@@
Step 3: Finally, we ask the question: is it possible that a non-zero no-crossing probability (that breaks uniform precession) explains the full Tsirelson violation? The first term on the right-hand side of Eq. 3.2 is positive, and we have calculated $\frac{1}{2}(1+\langle A\rangle)=-0.0598$ and $L_{1}=0.0421$ (from Appendix F). It is immediately clear that the violation of uniform precession is not sufficient to explain the Tsirelson inequality violation, hence non-zero quantum interferences are present. This is a crucial result, showing that this method has proven successful in identifying that uniform precession is satisfied well enough that quantum behaviour must be implied.
We also note that Eq. 3.2 can instead be written in terms of the two-time quasi-probabilities (Eq. 1.6) with no interferences:
@@
From Eq. 3.6 and Eq. 3.7, we see that the Tsirelson violation, $\frac{1}{2}(1+\langle A\rangle)$, will itself respect the Lüders bound if uniform precession is satisfied exactly ($L_{1}=0$), but in general can exceed this bound, as is the case with the maximally-violating state in the full infinite-dimensional QHO Hilbert space, which Tsirelson found to be $\frac{1}{2}(1+\langle A\rangle)=-0.13<-0.125$ [1].
@@
with the appropriate normalisation factor $N$, yields a Tsirelson violation of $\langle A\rangle=-1.0795<-1$. Thus, the method of assessing UP with an LG quantity can also be used upon such states, which may be of greater interest experimentally, for Gaussian state superpositions can be easier to create in a laboratory setting [72].
@@ III.3 Method 2: Three-time Sequential Measurements
Here $I$ is the interference term, and its values for each of the four cases are given below in Table 2. Both classically and quantum-mechanically, the triple correlator satisfies $|D|\leq 1$, which renders the first term $\frac{1}{2}(1-D)$ manifestly non-negative, so that Eq. 3.9 is the explicit form of Eq. 3.1 in this method.
Step 3: We compare the size of the UP violation, $2|\Delta p|$, with the size of the Tsirelson violation, $\frac{1}{2}|(1+\langle A\rangle)|$, noting once again that the interference terms would vanish in a purely classical model of the violation. Specifically, $2|\Delta p|=\frac{1}{2}|(1+\langle A\rangle)|$ marks the threshold that determines whether or not there is some plausible value of $D$ that can satisfy Eq. 3.9 with $I=0$.
Table 2 catalogues our numerical results with this approach, considering each of the four aforementioned distributions. We compare the extent of a lower-bound violation that would be possible from only the breaking of the dynamical assumption in a classical model with the true value of the Tsirelson violation, using the same state as before (Eq. 2.10). With the three-time sequential measurement probability $\Delta p_{123}$ as the chosen measure of UP, we calculate that $2|\Delta p_{123}|=0.0191<0.0598$, showing once again that the breaking of the dynamical assumption falls significantly short of accounting for the full extent of the Tsirelson violation, thereby implying the presence of quantum-mechanical interference terms.
When weak measurements are employed, UP is satisfied only marginally in the case of $\Delta p_{1(2)3}$; and when the quasi-probability distribution $\Delta q$ is instead used, which produces no interference terms, the UP violation is exactly commensurate with the Tsirelson violation. This is to be expected, since $\Delta q=0$ is a sufficient condition to derive the Tsirelson inequality, which property follows directly from the moment expansion form of Eq. 1.7. The quasi-probabilities, though they can be negative, respect the sum rules for probabilities; and when a measurement is rendered weak, the probability sum rule holds valid over that particular measurement time, as in $p_{1(2)3}(+,+,+)+p_{1(2)3}(+,-,+)=p_{13}(+,+)$.
@@
An alluring question is whether it is possible, by some variation of the parameters for the quantum state, to render $\Delta p_{123}=0$ exactly, whilst keeping within a Tsirelson violating region (such as that depicted in Fig. 4), which may serve the poetic justice of “total satisfaction” of the dynamical assumption, whereby the Tsirelson violation is solely due to effects quantum-mechanical in origin. In fact—unlike in Method 1, wherein setting $L_{1}=0$ leads trivially to the satisfaction of the Tsirelson inequality—it is possible in this case to set $\Delta p_{123}=0$. We find by using a constrained optimisation procedure via Lagrange multipliers that a simple state—still within the $N=6$ subspace, albeit now with complex coefficients—can be found to produce $\langle A\rangle=1.0756$ with $\Delta p_{123}=0$. The computational procedure is documented in Appendix J.
@@ IV.1 Dwell time
The dwell time is a measure of the fraction of time the particle spends in $x>0$. Classically, this value should be $1/2$ over a full period ($\tau=2\pi$). This can readily be turned into a self-adjoint operator, $\hat{T}_{D}(\tau)$ [73]. The average quantum-mechanical dwell time $\langle\hat{T}_{D}(\tau)\rangle$ can be ascertained experimentally by measuring $\langle\hat{\theta}(x(t))\rangle$ at multiple times.
@@
We see that the dwell time, integrated over a full period, is exactly the classical value. By this measure, a classical and quantum oscillator are indistinguishable.
@@
where $\tau$ is some time on the interval $0,2\pi]$, as the result was obtained using the mean value theorem. Thus, we see that the dependence of $\langle A\rangle$ upon the average dwell time drops out, since $\langle\hat{T}(2\pi)\rangle$ assumes the classical value, and we see that the Tsirelson violation is fundamentally rooted in the larger-than-classical probability current at the origin.
@@ IV.2 Crossing number
Another simple property of classical uniform precession, which is closely linked to Tsirelson’s dynamical assumption, is that trajectories cross the origin exactly twice over a period (barring the singular case of a particle sitting at rest at the phase-space origin). Classically, we can compute the number of crossings over a time interval $\tau$ using
@@
Explicit calculations indicate an asymptotic tendency towards the classical value of two crossings in the limit of high energy eigenstates, which is depicted in Fig. 6. Moreover, we calculate for the Zaw–Scarani state an average crossing number of $\langle\hat{N}_{c}(2\pi)\rangle_{Z}\approx 1.49$, with a standard deviation of $\Delta\hat{N}_{c}(2\pi)_{Z}\approx 0.2$. Over a third of a period, between measurement times, $\hat{N}_{c}$ remains diagonal for states in $\mathcal{H}_{0}$, and so $\langle\hat{N}_{c}(2\pi/3)\rangle_{Z}\approx 0.50$ with a standard deviation of $\Delta\hat{N}_{c}(2\pi/3)_{Z}\approx 0.07$. These results indicate that the probabilities of non-classical behaviour, such as two crossings during a third of a period, are very small indeed.
@@ IV.3 Implications for trajectory models
We have seen that the quantum-mechanical properties of the dwell time are identical to its classical properties and that as for the crossing number, its quantum-mechanical properties are qualitatively the same as classical and quantitatively similar. These features suggest that measurements of the dwell time and crossing number may be used to give indirect but useful bounds on the size of uniform precession violations. These quantities are, in principle, related to the three-time probabilities used in Section III, but any such relation will be non-unique, and we have not been able to find a particularly useful one. We proceed using a more heuristic analysis.
@@
The properties of the dwell time, Eq. 4.3, over the time interval $0,2\pi]$ indicate that the underlying trajectories must each spend equal amounts of time in $x>0$ and $x<0$. However, as noted, the dwell time must have the same properties over any shifted time interval of the same size. The implication in phase space is the trajectories must spend equal times above or below any infinite radial line. Thus, the set of trajectories must have rotational symmetry, which suggests that the only possibility is a set of circles around the origin with constant angular velocity along each circle. This still allows a set of circular trajectories with a spread of possible constant angular velocities, which will cause a violation of uniform precession. However, trajectories with angular velocities faster or slower than the classical one for the given Hamiltonian will create non-classical values for the crossing number. But, as we have seen above, the probabilities for such non-classical values of crossing numbers are very small.
@@ IV.4 Classical versus quantum dynamics
Finally, we make some general remarks about dynamics in classical versus quantum theories and why we might expect, on general grounds, that the violation of uniform precession be small. The predictions of quantum theory are very different to the predictions of classical physics, and most of these differences trace back to the simple fact that quantum theory allows a much richer variety of initial states than does classical physics—for example, through Wigner functions that can be negative versus classical phase-space probabilities. This fact is responsible for some of the most striking quantum phenomena, such as Bell inequality violations.
However, classical and quantum physics look at lot more similar at the level of dynamics, and there are many results in the history of quantum theory that prove this. For example, wave packets—or more generally, phase-space localised quasi-projectors—evolve to a good approximation along classical paths (see for example Ref. [74] and references therein). This in turn is related to the fact that the Wigner function, for a class of potentials, has approximately classical evolution. This behaviour can be affected by monitoring (e.g. with projection operators onto ranges of position at intermittent times), but this effect is not large unless the projections are onto very small regions and the projectors are applied at time intervals shorter than the energy time $\hbar/\langle E\rangle$ [75]. (For the Zaw–Scarani state considered in this paper, the energy time is calculated to be $1/12$ of a period.) Furthermore, in specific models, such as LG2 violations for a Gaussian intial state with harmonic oscillator dynamics, there is some non-classicality, but only about $30\%$ of the level of non-classicality that can be produced by a negative Wigner function [65, 66]. Of course, this discussion invokes quantum-mechanical properties, but the sceptical macrorealist could confirm these ideas using control experiments in which the dynamics of a system in a set of initially classical-like states are measured.
@@ V Summary and Discussion
But, as we noted in Section I, there is a crucial loophole in the Tsirelson inequality: particularly, that violations could ensue from the failure of the dynamical assumption of uniform precession in a classical system, meaning that a Tsirelson violation is not a sufficient condition for genuine quantum behaviour. Therefore, in this paper we have developed a subsidiary protocol to disentangle violations of macrorealism from the breaking of the dynamical assumption, to which end we have embedded the Tsirelson inequality within the wider Leggett–Garg formalism.
@@
In Section IV, by considering the heuristic notions of dwell times and the crossing number operator, we presented further qualitative evidence that uniform precession is satisfied well in states that violate the Tsirelson inequality. We found that the quantum-mechanical dwell time obeys the classical bound exactly and that the total number of origin crossings is approximately the classical value. On their own, these stand as interesting mathematical results, further elucidating the various ways in which uniform precession is and is not equivalent in the classical and quantum cases. But we have also further interpreted these quantities in the context of underlying trajectory models, noting that they greatly restrict the extent to which underlying trajectories can be erratic. We also came upon the global property that Tsirelson violations are linked to the probability currents through the origin, whereby weakly-flowing states, or the classicalised currents of quantum states, never produce a violation.
@@ Acknowledgements.
We are grateful to Clement Mawby, Shayan Majidy and Lin Htoo Zaw for useful discussions and comments on earlier versions of this manuscript, and the latter for bringing to our attention a result regarding superpositions of coherent states.
@@ Appendix A Tsirelson Quantity from the Wigner Function
Thus, The Tsirelson quantity can also be written in terms of the Wigner quasiprobability distribution $W_{\rho}(x,p)$ as
@@
where $W_{A}(x,p)$ (and likewise $W_{\rho}(x,p)$) is the phase-space representation of the operator $A$, obtained from its Weyl transform:
@@ E.1 Proof Strategy
    The largest matrix element, in magnitude, on a diagonal i.e for fixed $m$, $|Q_{n,n+2m+1}|$ is $|Q_{0,2m+1}|$.
@@
    The non-zero matrix elements decrease in magnitude across the 0th row i.e $|Q_{0,2m+1}|$ is a decreasing function in $m$.
@@ Appendix F Calculation of correlators in the quantum harmonic oscillator
$e^{i\hat{H}\tau}=\sum_{m=0}^{\infty}e^{im\tau}|m\rangle\langle m|\ ,$ (F5)
@@
$C_{12}=\frac{1}{2}\sum_{n,k=0}^{N}c_{n}^{*}c_{k}\langle n|\hat{Q}_{1}\sum_{m=0}^{\infty}e^{im\tau}|m\rangle\langle m|\hat{Q}_{1}e^{-ik\tau}|k\rangle+\frac{1}{2}\sum_{n,k=0}^{N}c_{n}^{*}c_{k}\langle n|e^{in\tau}\hat{Q}_{1}\sum_{m=0}^{\infty}e^{-im\tau}|m\rangle\langle m|\hat{Q}_{1}|k\rangle\ ,$ (F6)
@@
$\frac{1}{2}\sum_{n,k=0}^{N}\sum_{m=0}^{\infty}e^{i(m-k)\tau}c_{n}^{*}c_{k}Q_{nm}Q_{mk}+\frac{1}{2}\sum_{n,k=0}^{N}\sum_{m=0}^{\infty}e^{i(n-m)\tau}c_{n}^{*}c_{k}Q_{nm}Q_{mk}\ .$ (F7)
@@
$\frac{1}{2}\sum_{n,k=0}^{N}\sum_{m=0}^{\infty}e^{i(m-k)\tau}c_{n}^{*}c_{k}Q_{nm}Q_{mk}+\frac{1}{2}[\sum_{n,k=0}^{N}\sum_{m=0}^{\infty}e^{i(m-k)\tau}c_{n}^{*}c_{k}Q_{nm}Q_{mk}]^{*}\ .$ (F8)